40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Mon Jul 26 01:30:42 2021, Koji, Summary, Cryo vacuum chamber, Prep for the 2nd cooling of the suspension 
    Reply  Mon Jul 26 22:49:00 2021, Koji, Summary, Cryo vacuum chamber, Aquadag painting 10x
       Reply  Wed Jul 28 00:59:47 2021, Koji, Summary, Cryo vacuum chamber, The test mass successfully suspended P_20210727_154143.jpegP_20210727_190356.jpegP_20210727_190426.jpegP_20210727_190543.jpeg
          Reply  Thu Jul 29 00:42:38 2021, Koji, Summary, Cryo vacuum chamber, The test mass successfully suspended 
             Reply  Thu Jul 29 13:11:17 2021, Koji, Summary, Cryo vacuum chamber, Cooling progress: Update temp_log_cool_down_20210728_1830.pdfcooling_model1.pdfcooling_model2.pdfOSEM_cooling.pdf
                Reply  Fri Jul 30 12:22:56 2021, Koji, Summary, Cryo vacuum chamber, Cooling curve comparisons comparison_cold_head.pdfcomparison_inner_shield.pdfcomparison_test_mass.pdf
                   Reply  Sun Aug 1 22:22:00 2021, Koji, Summary, Cryo vacuum chamber, Cooling update temp_log_cool_down_20210728_1830.pdfcooling_meas.pdfOSEM_cooling.pdfcooldown_210728.zip
                      Reply  Mon Aug 2 13:25:32 2021, Stephen, DailyProgress, Cryo vacuum chamber, Warmup started 02 August 
                         Reply  Tue Aug 3 23:56:00 2021, Koji, DailyProgress, Cryo vacuum chamber, Warmup started 02 August P_20210804_000247.jpgP_20210803_235421.jpg
Message ID: 2621     Entry time: Thu Jul 29 00:42:38 2021     In reply to: 2620     Reply to this: 2622
Author: Koji 
Type: Summary 
Category: Cryo vacuum chamber 
Subject: The test mass successfully suspended 

[Stephen Koji]

Road to cooling down

  • The suspension with the test mass was installed in the chamber again
  • Looking at the oplev beam, we jiggled the wire loop position to adjust the alignment approximately.
  • The oplev beam was aligned more precisely.
     
  • We intentionally kept the OSEM at the "fully-open" position, while it is still close to the magnet so that we can have some actuation.
  • The coil driver was tested before closing the chamber, but it did not work.
    The coil itself was still intact, and the mirror was responding to the coil current if the coil current of ~100mA was applied from a bench power supply with the current ~100mA).
    So the problem was determined to be external.
     
  • Once we were satisfied with the oplev/OSEM conditions, the inner and outer lids were closed. Then the chamber was closed.
     
  •  Started pump down.
  • Started cooling down @18:30 / started temp logging too. Log filename: temp_log_cool_down_20210728_1830.txt

The photos were uploaded to Google Photo of WB labs.


The coil driver issue was resolved:

  • It was necessary to take care of the enable switch. Made a DB9 short plug for this purpose.
  • The output R was 1.2K (i.e. 2.4K across the + and - outputs). We needed ~10x more to see visible motion of the mass
  • e.g. The internal gain of the driver is x1.1. If we connect 5VDC input across the diff input of the driver yields, +11V shows up across the outputs of the final stage.
    If the R across the coil is ~100Ohm, we get ~100mA.
  • Soldered 6 x  330Ohm (1/8W) in parallel to 1.2K R_out. -> This ended up 51.5Ohm x2 across the coil. Each R=330 consumes ~1/10W. ->OK

Checking the DAQ setup / damping loop

  • DAQ setup
    • ADC: QPD X->FM16 / Y->FM17 / S->FM18 / OSEM-> FM19
    • DAC: CH11 -> Coil Driver In
  • Connected FM16 and FM17 to the coil drive by setting C4:TST-cdsMuxMatrix_12_17 and C4:TST-cdsMuxMatrix_12_18 to be 1.0
  • It was not obvious if the coil could damp the rigid body modes.
    • Actating the magnet caused Yaw motion most. Some Pitch motion too.
    • Configured FM16 and FM17 for the damping loop.
      • Filter Bank #1: [Diff0.1-10]  Zero 0.1Hz / Pole 10Hz
      • Filter Bank #10: [Anti Dewht]  Zero 1&200Hz / Pole 10&20Hz
    • Tried various damping gain. The mass was moving too much and the proper gain for the damping was not obvious.
    • So, the initial damping was obtained by shorting the coil at the coil in of the sat amp unit. (Induced current damping)
    • Once the test mas got quieter, it was found that -0.01 for FM16 could damp the yaw mode. Also it was found that +0.1 for FM17 could damp the pitch mode. (But not at once as the filters were not set properly)
       
  • TF measurement for calibration
    • The beam was aligned to the QPD
    • The test mass was damped by using the damping loops alternately 
    • Taken a swept sine measurement Filename: OSEM_TF_210729_0243.xml
      Recorded the time, saved the data, and took a screenshot
      • This measurement was taken @T_IS=252K / T_TM=268K @t=8hr (2:30AM), Rcoil=15.6Ohm
    • Second measurement Filename: OSEM_TF_210729_2147.xml
      • @T_IS=172K / T_TM = 201K @t=27.5hr (10PM), Rcoil=10Ohm
    • 3rd measurement Filename: OSEM_TF_210730_1733.xml
      • @T_IS=116K / T_TM = 161K @t=47hr (5:30PM), Rcoil=?
    • 4th measurement Filename: OSEM_TF_210731_2052.xml
      • @T_IS=72K / T_TM = 134K @t=75hr (9:30PM), Rcoil=6.0Ohm

OSEM LED/PD

  • The Satellite amp brought from the 40m is used as-is.
  • The initial OSEM reading was 8.8V, this corresponds to ~30000cnt.
  • As the OSEM was cooled, this number was increasing. To avoid the saturation, a voltage divider made of 4x 15kOhm was attached. I didn't expect to have the input impedance of the AA filter (10K each for the diff inputs), this voltage divider actually made 18.24V across POS and NEG output to be 5.212V to the AA fiter. So the voltage division gain is not 0.5 but 0.2859.
  • This made the ADC range saved, but we still have a risk of saturating the PD out. If this happens. The PD TIA gain will be reduced before warming up.
    -> The TIA and whitening stages use AD822, and the diff output stage uses AD8672. AD822 can drive almost close to rail-to-rail. AD8672 can drive upto ~+/-14V.

There was not enough time for the QPD calib -> Tomorrow

ELOG V3.1.3-